# Cloud gap-filling with deep learning for improved grassland monitoring

### v1

@@ 4.3.3 Model training
The CB includes an independent branch for each input layer. Each branch comprises two Conv1D layers with filter sizes of 8 and 16, both using a kernel size equal to 3. These are followed by a max pooling layer with a pool size of 3 (without striding), and finally, two fully-connected layers, with 32 and 16 units, respectively. To determine the optimal configuration, we conducted experiments with various parameter sets.
@@
The EDB was constructed using LSTM units with a 16-unit latent space. For the training, we employed the Adam optimizer, together with the Mean Squared Error (MSE) loss function. The latter, will allow us to effectively penalize significant errors that might arise from substantial discrepancies between predicted and target values, particularly when dealing with noisy data, such as preserved cloudy measurements. The initial learning rate was set to 0.005, and the batch size to 256. The training procedure included early stopping, which would terminate the process if the loss function did not improve after three consecutive epochs, with a maximum of 30 epochs to prevent overfitting.
@@
In addition, incorporating the NDVI time series as both input and target can result in the model being prone to overfitting to those specific values. To address this issue and force the network to give higher attention on masked timestamps, we have used temporal weighting for each input sample. Masked NDVI values were assigned a training weight parameter of $\mathrm{w_{\alpha}}$, while non-masked values were given a smaller weight parameter of $\mathrm{w_{\beta}}$. These parameters were set to be 0.75 and 0.25 respectively since they provided the optimal results. Additionally, as we utilized interpolation to produce a complete target time series, the values in timestamps that are not actual measurements and have been derived from the interpolation are assigned a weight of 0.
@@
Finally, to ensure an unbiased evaluation of both the NDVI reconstruction and the grassland event detection tasks, we selected the training set from randomly selected parcels across the study sites. For the NDVI reconstruction assessment, we also selected random pixels from parcels across the test sites. To avoid any bias in the subsequent evaluation of mowing events, we excluded all the pixels within the parcels that were annotated through the photo-interpretation process (Section 3.2.2) from the training set. Table 4 shows the numbers of parcels and pixels of the training and test sets in the six regions, for the NDVI reconstruction task.
@@
Table 4: Number of training and inference pixels and parcels for each region.
Region # training pixels (# parcels) # inference pixels (# parcels) Region 1 408,132 (2002) 86,747 (367) Region 2 184,452 (564) 140,242 (253) Region 3 68,180 (364) 36,715 (108) Region 4 158,419 (737) 47,785 (107) Region 5 49,377 (199) 25,502 (86) Region 6 114,742 (451) 28,307 (101) Total 983,302 (4317) 365,298 (1022)
@@ 4.4 Event detection
Mowing Detection Algorithm I (MDA I) is the algorithm developed by devroey2022 , in the context of Sen4CAP project. Its Sentinel-2 component of the processing chain for mowing detection searches for significant and sharp drops between consecutive NDVI cloud-free observations, surpassing a predefined threshold. This algorithm is considered a solid method for mowing event detection and similar methodologies have been widely used the past years. The evaluation of its performance across several European countries for 2019 demonstrated f1-score of 79% over 150 parcels in the Netherlands. However, when evaluated on 118 cases in Lithuania, it showed a notably lower f1-score of 60%, mainly due to a significantly reduced precision of 49%.
@@
Mowing Detection Algorithm II (MDA II) is the algorithm proposed by schwieder2021 . They suggest to consider the residuals of the assessed vegetation index values from an interpolated time series formulated based on the seasonal peak timestamps that represent the ideal growing season curve. This algorithm achieved an overall f1-score of 67% on 180 grassland parcels in Germany for 2020. The repository for their implementation was collected and cloned as a user-defined function in the FORCE processing environment [^1] https://github.com/davidfrantz/force-udf/tree/main/python/ts/mowingDetection.
@@
Both these methods, while not perfect, presented quite satisfactory results in our dataset. Nonetheless, we want to assess whether using SF to generate artificial cloud-free NDVI time series, we can improve the mowing event detection accuracy. Thus, we use these methods as baseline, since they are highly dependent on the NDVI information and can directly highlight the importance of the proposed methodology.
@@
Deep Learning Mowing Detection Algorithm (CNN-RNN): Apart from the fusion task, we utilize the SF architecture as an event detection algorithm. However, for the aim of this task, we keep only the NDVI branch at the input of the network since the purpose now is to evaluate the mowing detection accuracy given the full NDVI time series provided. Furthermore, we replaced the existing activation function at the end of the output layer with a sigmoid one, in order to quantitatively assess the probability of an event occurring at each distinct timestamp. For the training of the model we utilized the annotated labels that were derived through photo-interpretation, as mentioned in Section 3.2.2. For this purpose, we created binary series of 0s and 1s with a size equal to the length of the time series (in our case, 29). In this binary representation, the value 1 corresponds to the closest timestamp to the annotated mowing event. This labeling format allows for the model to learn intricate temporal patterns, enabling it to offer not just event detection but also a probability estimation for the likelihood of an event occurring at each timestamp within the extensive range of evaluated dates. This dual-purpose application of the SF architecture will showcase its versatility in addressing complex tasks beyond NDVI reconstruction, highlighting its potential as an effective event detection algorithm, too. It’s worth noting that while we acknowledge the potential contribution of SAR branches to the model’s accuracy, their inclusion in this model configuration would introduce an unfair advantage. Our objective here is to assess and quantify the unique contributions of the same optical features through different models, ensuring a fair and robust evaluation.
@@ 5 Results
In this section we evaluated the SF model’s performance in cloud gap-filling. We compared its performance with other common interpolation methods, namely Linear Interpolation (LI), Akima Interpolation (AI) and Quadratic Interpolation (QI). Additionally, we assessed the effect of missing NDVI images on the detection of grassland events and the ability of the SF model to alleviate this problem.
@@ 5.1 Gap-filling model performance
This section evaluates the performance of the SF model on the six grassland regions. For the evaluation we used the metrics of Mean Absolute Error (MAE) (Eq. 7) and coefficient of determination ($\mathrm{R^{2}}$) (Eq. 8), considering all available time steps with a cloud-free NDVI value.
@@
In Figure 7, we present scatter plots for each of the six study regions. Notably, we have noted a strong correlation between the ground truth and the predictions generated by the SF model. Region 5 stands out with relatively poorer performance, characterized by a MAE of 0.028 and a $\mathrm{R^{2}}$ value of 0.88. This discrepancy can be attributed to the presence of relatively high cloud coverage and a limited number of training pixels from the area. However, when assessing the other regions, we did not find any statistically significant differences in performance. Collectively, the $\mathrm{R^{2}}$ coefficient measures at 0.92, and the MAE is calculated at 0.024, which highlights on SF model’s capacity in effectively generating missing NDVI values.
@@
Fig. 8 presents the distribution of MAE for the SF model and the three alternative interpolation methods. This analysis focuses on timestamps that were intentionally masked in order to highlight the model’s ability to fill missing NDVI values. The SF model outperforms the temporal interpolation methods, with a mean MAE of 0.036 and a median of 0.032. In contrast, the AI and LI methods exhibit less favorable results, with average MAE values of 0.043 and 0.044, respectively. The QI method performs the poorest among the approaches, with an average MAE of 0.047 and a notably higher standard deviation.
@@
For further details regarding cloud gap-filling performance across various parameters such as study site, date of inference, cloud coverage, and length of consecutive missing values, please refer on A Figures S4 to S7.
@@ 5.2 Event detection accuracy assessment
In this section, we evaluate the SF methodology’s contribution to the detection of grassland mowing events. We use event timestamps as labels to assess the impact of the artificially created NDVI time series on the event detection task. We examined 803 unique grassland parcels across six regions, which correspond to 287,281 pixels, and derived from the photo-interpretation process (Section 3.2.2).
Our evaluation investigates how the SF model affects grassland event detection by comparing two rule-based methods outlined in Section 4.4. These methods depend exclusively on NDVI time series to identify substantial drops in their values. Consequently, they may exhibit limitations in detecting events, especially when consecutive cloudy observations are present. They provide a time range for each event, allowing us to compare it directly with the respective range obtained from photo-interpretation. Finally, we measure the detection accuracy and the ability to predict an event correctly using recall (Eq. 9) and precision metrics (Eq. 10), and their harmonic mean using the f1-score metric (Eq. 11). The predictions are aggregated and evaluated at the parcel level.
@@
A predicted event is deemed correct if its predicted date falls within twelve days (tolerance parameter) before or after the actual event, which is twice the size of the time series interval. Events falling outside this tolerance window are considered false positives.This approach helps mitigate issues arising from the 6-day interpolation used for detection. This may cause events to shift from one timestamp to another, introducing challenges when reporting results, as it can impact the accuracy and alignment of detected events with their actual occurrences. Furthermore, while MAE is a common metric for evaluating mowing event detection algorithms, it is not considered applicable in our case since we lack the exact date of the mowing events. Our main purpose in this study is to assess how the implementation of the SF model for reconstructing missing NDVI values enhances mowing event detection by identifying cases that would have been missed otherwise. Nevertheless, in A (Tables S1-S3), we present the respective results for various evaluation approaches, which highlights the independence of the SF model on the criteria used to define a predicted event as correct.
Table 5 shows the results of MDA I and MDA II algorithms acquired from the different interpolation methods and time series without any interpolation applied. Both algorithms demonstrate similar performance, with slight variations depending on the interpolation methodology. As indicated in the table, all interpolation approaches exhibit a common tendency to slightly underestimate recall while significantly enhancing precision accuracy for both mowing detection algorithms, compared to using unprocessed time series data. This can be attributed on the interpolation capabilities to effectively eliminate numerous abrupt changes that often arise from noisy NDVI successive values by providing consistent timestamps with a 6-day gap between them. All in all, this behaviour decreases the total number of false detected mowing events (i.e., increased precision), reducing the number of some correctly identified cases though (i.e., reduced recall). In essence, these findings highlight the compromise between precision and recall, emphasizing the role of interpolation methods in refining mowing event detection accuracy. Specifically, the SF provided NDVI time series, present the best results in both MDA I and MDA II, with the latter achieving the optimal f1-score equal to 0.84. For MDA I, recall presents lower values for all interpolation methods comparing the raw NDVI time series, while for MDA II, SF managed even to improve detection accuracy by almost 1% compared to raw time series. Furthermore, SF offers significantly higher precision (over 90%) for both MDA I and II, with a range of 19% to 24% improvement over the initial NDVI time series.
@@
In Fig 9, we present f1-score categorized by cloud coverage percentage (ccp). We divided the groups using the mean ($\mathrm{\mu}\simeq 0.4$) and standard deviation ($\mathrm{\sigma}\simeq 0.1$) values of the cloud coverage distribution, which resembled a normal distribution. Based on this statistical analysis, we constructed four distinct categories accordingly: ccp $\mathrm{<\mu-\sigma}$ (101 cases), $\mathrm{\mu-\sigma\leq}$ ccp $\mathrm{<\mu}$ (252 cases), $\mathrm{\mu\leq}$ ccp $\mathrm{<\mu+\sigma}$ (291 cases), ccp $\mathrm{\geq\mu+\sigma}$ (159 cases). The figure reveals that SF’s delivers consistently superior performance across the various cloud coverage scenarios, which highlights its robustness against any cloud coverage situation. This is becoming more evident in cases with very high cloud coverage percentage. This can be attributed to SF’s ability to handle cases resembling extended cloud coverage, which often obscure mowing events. The NDVI time series constructed based on SAR data excels in recovering such challenging cases, resulting in better performance.
Figure 9: F1-score for mowing event detection and for different cloud coverage percentage (ccp) scenarios. For further analysis, these results are detailed in Table S5 in A.
Fig. 10 illustrates the recall, precision, and f1-score obtained from both MDA I and MDA II across each of the study regions. Once again, the results obtained from MDA II exhibit a slight improvement over MDA I. Notably, Region 3 stands out with optimal results with precision values close to 1 and very high recall rates, which results to f1-score surpassing 0.9 for both MDAs. Similar behavior exhibit Regions 4 and 5, too. On the other hand, the rest of the Regions present worse performance, with and f1-score around 0.80. It is important to highlight that overall, precision can reach significantly high values, occasionally approaching 1, when we utilize the NDVI time series constructed with the SF model. This effect can be attributed to SF model’s ability to correct cloudy observations and thus prevent the MDAs to incorrectly identify these cases as mowing events.
@@
In Fig. 11, a scatter plot is presented that compares the ground truth and predicted values using MDA II, which has proved to be slightly more robust that MDA I. Only absolute differences less than 12 days are considered accurate, as larger gaps are deemed too significant. The size of each point indicates the number of observations on this reference day. The plot shows a linear relationship between the measures, with an $\mathrm{R^{2}}$ value of 0.946, and most events occurring between June and July. Misclassifications mainly arise when the model detects events slightly later than their actual occurrence, possibly due to delayed NDVI decreases resulting from prior missing values.
Figure 11: Scatter plot comparing the grassland event reference Day of Year (DoY) and predicted DoY using MDA II for event prediction and SF algorithm for interpolation of the NDVI input time series. The size of each point indicates the number of observations.
Fig. 12 illustrates the recall for MDA II, for the different number of mowing events annotated per parcel. Among the 54 unmown parcels, approximately 79.6% were correctly identified as unmown. Furthermore, in parcels with a single mowing event, the recall percentage demonstrates a relatively improved performance, surpassing 85%. However, in the case of the 95 parcels with two mowing events, only in 29.5% of them the algoirthm managed to accurately identified both events. Remarkably, in approximately 76% of these two-event parcels, at least one event was detected accurately.
Figure 12: Evaluation of the performance of MDA II on the different number of mowing events that occurred in parcels during the season. The first bar corresponds to the performance against non-mowed parcels, the second bar against parcels that were mowed exactly one time and the third bar for parcels that were mowed twice. For each of them the percentage of actual cases correctly identified (i.e., recall) or the portion of the actual events omission is illustrated.

### v2

@@ 4.3.3 Model training
The CB includes an independent branch for each input layer. Each branch comprises two Conv1D layers with filter sizes of 8 and 16, both using a kernel size equal to 3. These are followed by a max pooling layer with a pool size of 3 (without striding), and finally, two fully connected layers, with 32 and 16 units, respectively. The EDB was constructed using LSTM units with a 16-unit latent space. To determine the optimal configuration, we conducted experiments with various parameter sets. Ultimately, we settled on a lightweight configuration with a total of 82,337 parameters.
For the training, we employed the Adam optimizer, due to its adaptive learning rates and computationally efficient handling of non-stationary data, together with the Mean Squared Error (MSE) loss function. The MSE loss function is defined as:
$MSE=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (7)
@@
where $y_{i}$ represents the true values and $\hat{y}_{i}$ represents the predicted values. MSE is well-suited for our task because it penalizes large errors effectively that might arise from substantial discrepancies between predicted and target values. This is particularly beneficial for our scenario where the data can be noisy, containing preserved cloudy measurements. The initial learning rate was set to 0.005, and the batch size to 256. The training procedure included early stopping, which would terminate the process if the loss function did not improve after three consecutive epochs, with a maximum of 30 epochs to prevent overfitting.
@@
In addition, incorporating the NDVI time series as both input and target can result in the model being prone to overfitting to those specific values [28]. To address this issue and force the network to give more attention to masked timestamps, we have used temporal weighting for each input sample. Masked NDVI values were assigned a training weight parameter of $\mathrm{w_{\alpha}}$, while non-masked values were given a smaller weight parameter of $\mathrm{w_{\beta}}$. These parameters were set to be 0.75 and 0.25 respectively, since they provided the optimal results. Additionally, as we utilized Akima interpolation to produce a complete target time series, the values in timestamps that are not actual measurements and have been derived from the interpolation were assigned a weight of 0. This ensured they were not considered during the training and validation processes.
@@
Finally, for NDVI reconstruction, we selected the training and validation pixel sets from the complete parcel time series after the artificial masking process (see Section 4.2). To avoid any bias in the subsequent evaluation of mowing event detection, we excluded from the training set all the pixels within the parcels that were annotated through the photo-interpretation (Section 3.2.2). Table 4 shows the numbers of parcels and pixels of the training and validation sets in the six regions, for the NDVI reconstruction task.
@@
Table 4: Number of training and validation pixels and parcels for each region.
Region # training pixels
(# parcels)
# validation pixels
(# parcels)
Region 1 (Alytus) 408,132 (2,002) 86,747 (367) Region 2 (Klaipėda) 184,452 (564) 140,242 (253) Region 3 (Utena) 68,180 (364) 36,715 (108) Region 4 (Telšiai) 158,419 (737) 47,785 (107) Region 5 (Panevėžys) 49,377 (199) 25,502 (86) Region 6 (Šiauliai) 114,742 (451) 28,307 (101) Total 983,302 (4,317) 365,298 (1,022)
@@ 4.4 Event detection
We employed two distinct algorithms for mowing event detection, referenced as follows:
  1. 1.
    Mowing Detection Algorithm I (MDA I): This algorithm, developed by [55], was implemented using pseudo-code from the Sen4CAP package description [69].
  2. 2.
@@
    Mowing Detection Algorithm II (MDA II) is the algorithm proposed by [56]. The repository for their implementation was collected and cloned as a user-defined function in FORCE processing environment [^1] https://github.com/davidfrantz/force-udf/tree/main/python/ts/mowingDetection.
@@
Both these methods, while not perfect, presented quite satisfactory results in our dataset. Nonetheless, we want to assess whether using SF to generate an artificial cloud-free NDVI time series, can improve the mowing event detection accuracy. Thus, we used these methods as a baseline, since they are highly dependent on the NDVI information and can directly highlight the importance of the proposed methodology.
@@
Deep Learning Mowing Detection Algorithm (CNN-RNN): Apart from the fusion task, we utilized the SF architecture as an event detection algorithm. However, for the aim of this task, we kept only the NDVI branch at the input of the network, since the purpose was to evaluate the mowing detection accuracy given the complete NDVI time series provided. Furthermore, we replaced the existing activation function at the end of the output layer with a sigmoid one, to quantitatively assess the probability of an event occurring at each distinct timestamp. For the training of the model, we utilized the annotated labels that were derived through photo-interpretation, as mentioned in Section 3.2.2. For this purpose, we created binary series of 0s and 1s with a size equal to the length of the time series (in our case, 29). In this binary representation, the value 1 corresponds to the closest timestamp to the annotated mowing event. This labeling format allowed the model to learn intricate temporal patterns, enabling it to offer not just an event detection but also a probability estimation for the likelihood of an event occurring at each timestamp within the extensive range of evaluated dates. This dual-purpose application of the SF architecture showcased its versatility in addressing complex tasks beyond NDVI reconstruction, highlighting its potential as an effective event detection algorithm, too. It’s worth noting that while we acknowledge the potential contribution of SAR branches to the model’s accuracy, their inclusion in this model configuration would introduce an unfair advantage. Overall, our objective here is to assess and quantify the unique contributions of the same optical features through different models, ensuring a fair and robust evaluation.
@@ 5 Results
In this section, we evaluated the SF model’s performance in cloud gap-filling. We compared its performance with other common interpolation methods, namely Linear Interpolation (LI), Akima Interpolation (AI), and Quadratic Interpolation (QI) [68, 70]. More specifically, LI estimates values between known data points by drawing straight lines, AI uses piecewise cubic polynomials to smooth the curves between points, while QI fits quadratic curves between adjacent points, providing an even smoother interpolation compared to the rest. Additionally, we assessed the effect of missing NDVI images on the detection of grassland events and the ability of the SF model to alleviate this problem.
@@ 5.1 Gap-filling model performance
This section evaluates the performance of the SF model on the six grassland regions. For the evaluation, we used the metrics of Mean Absolute Error (MAE) (Eq. 8) and coefficient of determination ($\mathrm{R^{2}}$) (Eq. 9), considering all available time steps with a cloud-free NDVI value.
@@
In Fig. 7, we present scatter plots for each of the six study regions. Notably, we have noted a strong correlation between the ground truth and the predictions generated by the SF model. Region 5 stands out with relatively poorer performance, characterized by a MAE of 0.028 and a $\mathrm{R^{2}}$ value of 0.88. This discrepancy can be attributed to the presence of relatively high cloud coverage and a limited number of training pixels from the area. However, when assessing the other regions, we did not find any statistically significant differences in performance. Collectively, the $\mathrm{R^{2}}$ coefficient measures at 0.92, and the MAE is calculated at 0.024, which highlights on SF model’s capacity in effectively generating missing NDVI values.
@@
Fig. 8 presents the distribution of MAE for the SF model and the three alternative interpolation methods. This analysis focuses on timestamps that were intentionally masked in order to highlight the model’s ability to fill missing NDVI values. The SF model outperformed the temporal interpolation methods, with a mean MAE of 0.036 and a median of 0.032. In contrast, the AI and LI methods exhibited less favorable results, with average MAE values of 0.043 and 0.044, respectively. The QI method performed the poorest among the approaches, with an average MAE of 0.047 and a notably higher standard deviation.
@@
For further details regarding cloud gap-filling performance across various parameters such as study region, date of inference, cloud coverage, and length of consecutive missing values, please refer to A Fig. S4 to S7.
@@ 5.2 Event detection accuracy assessment
In this section, we evaluate the SF methodology’s contribution to the detection of grassland mowing events. We used event timestamps as labels to assess the impact of the artificially created NDVI time series on the event detection task. We examined 803 unique grassland parcels across the six study regions, which correspond to 287,281 pixels, and derived from the photo-interpretation process (Section 3.2.2).
Our evaluation investigates how the SF model affects grassland event detection by comparing two rule-based methods outlined in Section 4.4. These methods depend exclusively on NDVI time series to identify substantial drops in their values. Consequently, they may exhibit limitations in detecting events, especially when consecutive cloudy observations are present. They provide a time range for each event, allowing us to compare it directly with the respective range obtained from photo-interpretation. Finally, we measured the detection accuracy and the ability to predict an event correctly using recall (Eq. 10) and precision metrics (Eq. 11), and their harmonic mean using the F1-score metric (Eq. 12). The predictions are aggregated and evaluated at the parcel level of photo-interpretation.
@@
A predicted event is deemed correct if its predicted date falls within twelve days (tolerance parameter) before or after the actual event, which is twice the size of the time series interval. Events falling outside this tolerance window are considered false positives. This approach helps mitigate issues arising from the 6-day interpolation used for detection. This may cause events to shift from one timestamp to another, introducing challenges when reporting results, as it can impact the accuracy and alignment of detected events with their actual occurrences. Furthermore, while MAE is a common metric for evaluating mowing event detection algorithms, it is not considered applicable in our case since we lack the exact date of the mowing events. Our main purpose in this study is to assess how the implementation of the SF model for reconstructing missing NDVI values enhances mowing event detection by identifying cases that would have been missed otherwise. Nevertheless, in A (Tables S1-S3), we present the respective results for various evaluation approaches, which highlights the independence of the SF model on the alternative criteria used to define a predicted event as correct.
Table 5 shows the results of MDA I and MDA II algorithms acquired from the different interpolation methods and time series without any interpolation applied. Both algorithms demonstrated similar performance, with slight variations depending on the interpolation methodology. As indicated in the respective table, all interpolation approaches exhibited a common tendency to slightly reduce recall while significantly enhancing precision accuracy for both mowing detection algorithms, compared to using unprocessed time series data. This can be attributed on (i) the potential events omission due to optical data interpolation to Sentinel-1 intervals [59]; and (ii) the interpolation capabilities to effectively eliminate numerous abrupt changes that often arise from noisy successive NDVI values by providing consistent timestamps with a 6-day gap between them. All in all, this behavior decreases the total number of falsely detected mowing events (i.e., increased precision), reducing the number of some correctly identified cases (i.e., reduced recall). In essence, these findings highlight the compromise between precision and recall, emphasizing the role of interpolation methods in refining mowing event detection accuracy. Notably, the SF generated NDVI time series, presented the best results in both MDA I and MDA II, with the latter achieving the optimal F1-score equal to 0.84. Eventually, it appears that the SF methodology has effectively addressed challenges associated with the inference on Sentinel-1 intervals. More specifically, for MDA I, recall values are significantly better among the interpolation methods, while for MDA II, they are improved by almost 1% with regard to the raw time series (i.e., no interpolation applied). Furthermore, SF offered significantly higher precision (over 90%) for both MDA I and II, with a range of 19% to 24% improvement over the unprocessed NDVI time series.
@@
In Fig 9, we present F1-score categorized by cloud coverage percentage (ccp). We divided the groups using the mean ($\mathrm{\mu}\simeq 0.4$) and standard deviation ($\mathrm{\sigma}\simeq 0.1$) values of the cloud coverage distribution, which resembled a normal distribution. Based on this statistical analysis, we constructed four distinct categories accordingly: ccp $\mathrm{<\mu-\sigma}$ (101 cases), $\mathrm{\mu-\sigma\leq}$ ccp $\mathrm{<\mu}$ (252 cases), $\mathrm{\mu\leq}$ ccp $\mathrm{<\mu+\sigma}$ (291 cases), ccp $\mathrm{\geq\mu+\sigma}$ (159 cases). The figure reveals that SF consistently delivered superior performance across various cloud coverage scenarios, highlighting its robustness against any cloud coverage situation. This is becoming more evident in cases with very high cloud coverage percentages. This can be attributed to SF’s ability to handle cases resembling extended cloud coverage, which often obscure mowing events. The NDVI time series constructed based on SAR data excels in recovering such challenging cases, resulting in better performance.
Figure 9: F1-score for mowing event detection and for different cloud coverage percentage (ccp) scenarios. For further analysis, the results are detailed in Table S7 in A.
Fig. 10 illustrates the recall, precision, and F1-score obtained from both MDA I and MDA II across each of the study regions. Once again, the results obtained from MDA II exhibit a slight improvement over MDA I. Notably, Region 3 stands out with optimal results with precision values close to 1 and very high recall rates, which results in F1-score surpassing 0.9 for both MDAs. Similar behavior is exhibited in Regions 4 and 5, too. On the other hand, the rest of the Regions present worse performance, with an F1-score around 0.80. It is important to highlight that overall, precision can reach significantly high values, occasionally approaching 1, when we utilize the NDVI time series constructed with the SF model. This effect can be attributed to the SF model’s ability to correct cloudy observations and thus prevent the MDAs from incorrectly identifying these cases as mowing events.
@@
Fig. 11 depicts the recall, precision, and F1-score obtained from both MDA I and MDA II for NDVI differences between successive timestamps. Detection accuracy improves with larger NDVI differences, reaching near-perfect results for differences with absolute values greater than 0.2. However, for smaller drops (less than 0.1), detection becomes more challenging, particularly for MDA I, which identifies slightly more than half of the actual cases (recall of 55%). This issue does not necessarily reflect the quality of the NDVI time series produced by the model, but rather the sensitivity of the mowing detection algorithms to minor NDVI changes. As anticipated, due to the design of the applied methodologies, their efficiency decreases with lower NDVI differences, underscoring the need for adaptive thresholds.
Figure 11: Mowing detection performance (i.e., recall, precision and F1-score) across NDVI differences between successive timestamps for MDA I and MDA II, using NDVI time series produced by the proposed SF model.
In Fig. 12, a scatter plot is presented that compares the ground truth and predicted values using MDA II, which has been proven to be slightly more accurate than MDA I. Only absolute differences less than 12 days are considered accurate, as larger values are deemed too significant or distant. The size of each point indicates the number of observations on this reference day. The plot indicates a linear relationship between the measures, with an $\mathrm{R^{2}}$ value of 0.946. Most events occur in June and July, as indicated by the top bar plot. Misclassifications mainly arise when the model detects events slightly later than their actual occurrence, possibly due to delayed NDVI decreases resulting from prior missing values. The bottom bar chart presents the F1-scores grouped by month, revealing small fluctuations. This consistency in F1-scores across months highlights the model’s robustness in event timing prediction.
Figure 12: Scatter plot comparing reference and predicted Day of Year (DoY) for grassland events using MDA II. Point size indicates the number of observations. The top bar chart shows the number of events and the bottom bar chart shows the F1-Score per month.
Fig. 13 illustrates the recall for MDA II across parcels annotated with different numbers of total mowing events. Among the 54 unmown parcels, approximately 79.6% were correctly identified as unmown. Furthermore, in parcels with a single mowing event, the recall percentage showed significantly improved performance, exceeding 85%. However, for the 95 parcels with two mowing events, the algorithm accurately identified both events in only 29.5% of cases. In particular, the recall rate for the first mowing event was 66%, whereas for the second event, the recall rate dropped to 39%. This discrepancy could be explained if we combine Fig. 11 and Fig. S9. Second mowing event cases presented much lower absolute values of NDVI drops (lower than 0.1), where the respective performance of MDA II is worse. Notably, in approximately 76% of these two-event parcels, at least one event was detected accurately. Similar results for parcels with two mowing events have also been reported by [55] across Wallonia.
Figure 13: Evaluation of the performance of MDA II on the different number of mowing events that occurred in parcels during the season. The first bar corresponds to the performance against non-mowed parcels, the second bar against parcels that were mowed exactly one time and the third bar for parcels that were mowed twice. For each of them, the percentage of actual cases correctly identified (i.e., recall) or the portion of the actual events omission is illustrated.